calculate the volume of an ingot 40x50x400 mm long. it is cast into a cylinder 210mm diamter, calculate the lenght and surface area of this cylinder?
Answer by oberobic(2304)

Setup your variables:
length = l = 40
height = h = 50
depth = d - 400
Volume of Ingot = length * height * depth
V = 40 * 50 * 400 = 2000 * 400 = 800,000
...
Hint: As soon as I saw this big numbers, I started an Excel workbook to do my calculating. If you have a calculator, now would be a good time to go get it.
...
The volume of whatever is shaped into the ingot is to be cast into a cylinder 210 in diameter.
...
We know several more useful facts:
Volume of cylinder = area of base * height
Area = pi* radius^2
Radius = diameter/2 = 210/2 = 105
Area = pi * 105^2
...
So our unknown is the height of the cylinder.
But we know it has to have volume 800,000 to hold the same volume as the ingot.
800,000 = pi* 105^2 * height of cylinder
Solving for the height, we find:
ht = 23.097316
...
We can check our work by recomputing the volume and comparing it to the given volume.
Does pi*(105^2)*23.097316 = 800,000? Yes.
The height of the cylinder will be 23.097316 to have volume equal to an ingot 40 by 50 by 400.
...
Are we done yet?
...
Well, checking back to the question, we also are asked the surface are of the cylinder.
That is simply the circumference * ht of the cylinder.
ht = 23.097316
Circumference is just pi * diameter
C = pi * 210
C = 659.7344573
Surface area = C * ht
S = 659.7344573 * 23.097316
S = 15238.09524
...
Now we're done.